OpenStudy (anonymous):

solve each logarithmic equation. Express all solution in exact form. log base 2 (x-7)+ log base 2 X= 3

OpenStudy (anonymous):

\[\log_{n}(x)=\ln(x)/\ln(n)\]

OpenStudy (anonymous):

\[e^(\ln(x)=x \]

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

still kinda confused

OpenStudy (anonymous):

it become \[(\ln(x-7)+\ln(x))/\ln(10)= 3\]

OpenStudy (anonymous):

and you have also \[\ln(x-7)+\ln(x)=\ln(x*(x-7))\]

OpenStudy (anonymous):

where is the base of 2 going?